The length of a certain conductor of resistance 100 Ω is doubled and its cross-sectional area is halved. Its new resistance is


1
Expert's answer
2021-08-23T12:40:19-0400

The resistance of a condacting wire

"R=\rho\frac{l}{A}"

Hence,

"R'=\rho\frac{l'}{A'}=\rho\frac{2l}{A/2}=4\rho\frac{l}{A}=4R\\
=4*200\:\Omega=800\:\Omega"